# A comprehensive framework for multi-fidelity surrogate modeling with noisy data: a gray-box perspective

### v1

@@ 2.4 Implementation
2.4.1 Construction of a multi-fidelity gray-box model
Starting from a set of high-fidelity experimental data $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$ and a lower-fidelity computational model ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$, we provide here the methodology to combine the two to construct a gray box as a MFSM.
Our approach to MF gray-box modeling uses polynomial chaos expansion as a surrogate model in the hybrid correction scheme introduced in Equation 9, extending the works of Ng and Eldred (2012); Palar et al. (2016); Berchier (2016). The main motivations behind our choice of PCE as a surrogate in our MFSM methodology include its robustness to noise, its efficiency in terms of training, and its applicability in uncertainty quantification problems.
@@ 2.4.1 Construction of a multi-fidelity gray-box model
where $c_{\boldsymbol{\alpha}}\in\mathbb{R}$ are the coefficients of the multivariate polynomials $\Psi_{\boldsymbol{\alpha}},\,\boldsymbol{\alpha}\in{\mathcal{A}}$}. Each polynomial $\Psi_{\boldsymbol{\alpha}}$ is the product of univariate polynomials orthogonal with respect to the PDF $f_{\boldsymbol{X}_{i}}$ of the input variable $\boldsymbol{X}_{i}$, and characterized by the multi-index $\boldsymbol{\alpha}$. ${\mathcal{A}}\subset\mathbb{N}^{M}$ is the finite set of multi-indices of the polynomials, and it can be obtained from different truncation schemes, such as total-degree, low-rank or hyperbolic truncation (Marelli et al., 2022).
For the calculation of the PCE coefficients, we adopt a regression-based strategy, as exhaustively reviewed in Lüthen et al. (2021, 2022), because of its applicability to data-driven problems and robustness to noise (Torre et al., 2019). Specifically, we opt for a sparse regression solver, least angle regression (LARS) (Blatman and Sudret, 2011). Moreover, for the choice of the PCE basis we use degree adaptivity, as well as a total-degree truncation scheme for low-dimensional applications and hyperbolic truncation for higher-dimensional applications.
The first step in constructing our MF gray-box model as in Equation 9 entails creating a surrogate of the low-fidelity model using PCE. For this purpose, we first sample $N_{\text{L}}$ realizations $\boldsymbol{{\mathcal{X}}}_{\text{L}}=\{\boldsymbol{x}_{\text{L}}^{(1)},...,\boldsymbol{x}_{\text{L}}^{(N_{\text{L}})}\}$ of the input random variables (e.g. through Latin hypercube sampling; McKay et al. (1979)) and obtain the corresponding model responses ${\mathcal{Y}}_{\text{L}}=\{y_{\text{L}}^{(1)},...,y_{\text{L}}^{(N_{\text{L}})}\}$. We then construct a PCE model
@@
as discussed above. By using ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$ in place of the original LF model ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$, we eliminate the need for the HF training set to be a subset of the LF one, as from now on we are able to obtain evaluations of ^$\widehat{{\mathcal{M}}}_{\text{L}}$ at a negligible cost. Moreover, we are able to remove any noise that may be present in the LF data in a general MFSM scenario.
@@
where $\mathcal{A}_{\text{L}}\cap\mathcal{A}_{\delta}$ is the set of multi-indices present in both the LF and the discrepancy expansions, $\mathcal{A}_{\text{L}}\text{\textbackslash}\mathcal{A}_{\delta}$ is the set of multi-indices present only in the LF expansion, and $\mathcal{A}_{\delta}\text{\textbackslash}\mathcal{A}_{\text{L}}$ the set of multi-indices present only in the discrepancy expansion. This last step is optional, and it is only possible when both ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$ and $\hat{\delta}(\boldsymbol{x})$ are PCEs. In general, different SMs could be used for either of these two models, in which case the combined expression in Equation 19 is not available. Thus, the corresponding step is omitted, and the MF gray-box predictor is instead expressed as:
@@
In summary, the construction of the multi-fidelity gray-box model from a HF experimental data set $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$ and a LF computational model ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$ according to Equation 9 involves the following main steps:
@@
    Use the computed ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x}),\,\hat{\rho},\,\hat{\delta}(\boldsymbol{x})$ for the MF gray-box predictor as in Equation 20. Optionally, merge the LF and discrepancy expansions into one PCE.
In a broader multi-fidelity setting, a low-fidelity dataset, or even a pre-trained surrogate of the LF model (not necessarily a PCE model) can be available instead of a LF computational model. In this situation, one can follow the same procedure to construct the MFSM, by simply omitting Step $1$, and in the second case also Step $2$. Furthermore, when a high-fidelity computational model is available instead of a HF dataset, an additional step precedes Step $1$. This consists in using sampling to obtain a HF experimental design $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$. These adaptations accommodate the general multi-fidelity case, which need not strictly adhere to the grey-box modeling paradigm.
@@
one can jointly estimate $\rho$ and the coefficients $c_{\alpha,\delta,\text{new}}$ using OLS.
Another approach is based on alternating least squares (Chevreuil et al., 2015), where the estimates for $\rho$ and $\delta(\boldsymbol{x})$ are iteratively refined. The starting value for $\hat{\rho}$ in this joint optimization can be obtained from the 4th step of the algorithm described above.
In the applications discussed in Section 3, we opt for the method outlined in Steps 1-6 due to its simplicity. Notably, its performance closely paralleled that of the two variations discussed above.
@@ 2.4.2 Construction of confidence and prediction intervals
One of the applications of the bootstrap estimator lies in constructing confidence intervals for regression models (Freedman, 1981). Although the application of bootstrap to provide local error estimates to PCE model predictions within a single-fidelity context has been previously studied (see Marelli and Sudret (2018)), its usage in the context of MFSM has not yet been explored.
@@
This means that a realization of $\varepsilon_{\text{H}}$ can be estimated as the difference between a HF observation and the bootstrap mean. Having $N_{\text{H}}$ noise samples $r^{(i)}$, we use maximum likelihood estimation (MLE) to infer the parameters of a zero-mean distribution. Within the scope of this work, we consider zero-mean variants of the classical Gaussian, Laplace, and Uniform distributions, but in the general case any distribution could be considered. Finally, we use the Bayesian information criterion (BIC; Schwartz (1978)) to select the most appropriate distribution among those considered.
@@ 3 Validation and results
In this section, the performance of the proposed framework for multi-fidelity gray-box modeling is illustrated on three examples of increasing complexity: an analytical example with one-dimensional input (Section 3.3), a case study with a ten-dimensional finite-element model of a truss and its simply supported beam approximation (Section 3.4), and a real-world application involving wind turbine simulations (Section 3.5). In each application, the HF data contain noise, which is either naturally present (Example 3.5) or artificially introduced by us to replicate the gray-box scenario (Examples 3.3, 3.4). The validation of the proposed framework comprises two parts: assessing the performance of the MFSM constructed as described in Section 2.4.1, and appraising the confidence and prediction intervals constructed as outlined in Section 2.4.2.
@@ 3.1 MFSM performance and convergence evaluation
where $y_{\text{t}}^{(i)}$ equals the noise-free HF model response $\psi_{\text{H}}(\boldsymbol{x}_{\text{t}}^{(i)})$ at the test point $\boldsymbol{x}_{\text{t}}^{(i)}$, when this response is available (Examples 3.3, 3.4), or the noisy HF response $y_{\text{H}}(\boldsymbol{x}_{\text{t}}^{(i)})$ when the noise-free response is not known (Example 3.5), and $\mu_{y}$ is the mean value of the HF response.
@@ 3.3 Analytical 1-D example
where $x\sim\mathcal{U}\left[0,2\right]$. The HF and LF functions are depicted in Figure 2.
To replicate the MF gray-box setting, we artificially contaminate the HF data with additive noise that follows a Gaussian distribution $\varepsilon_{\text{H}}\sim\mathcal{N}(0,\,\sigma_{\varepsilon_{\text{H}}})$.
@@ 3.3.1 MFSM performance and convergence
Here, for each noise level, the HF ED varies from 5 to 25 data points, while the LF ED is fixed in all experiments to 100 data points. For the computation of the validation errors shown in LABEL:fig:1-d_boxplot, we use $N_{test}=10^{5}$ noise-free HF data points generated using LHS in the input space.
@@
We observe that from as few as 5 HF training data points, our MFSM approach outperforms surrogate models trained solely on either HF or LF data. This distinction becomes particularly evident for lower noise levels within the HF data. Hence, when HF data is scarce and data of different fidelities is present, the value of employing MFSMs as opposed to single-fidelity SMs becomes apparent. We notice that when sufficient HF data is available, e.g., here, approximately 20 data points for $\varepsilon_{\text{H}}\sim{\mathcal{N}}(0,\,0.01\hat{\sigma}_{\text{H}})$, the MFSM and HF surrogate model performance is similar. In addition, as the level of noise in the HF data increases, the difference in performance of the MFSM and the HF SM diminishes. Indeed, LABEL:fig:noise_4 shows a comparable performance between the two models, regardless of the size of the experimental design.
@@ 3.3.2 Confidence and prediction intervals
LABEL:fig:1d_CIs_PIs shows the $90\%$ CIs (blue area) and $90\%$ PIs (yellow area) for the MFSM prediction (blue line) in four different cases occurring from all combinations among a realization of higher/lower noise in the HF observations and a realization of a larger/smaller HF ED. Here, again the LF ED is fixed to 100 samples. The HF and LF training data in each case is visualised by the black error bars and the gray circles, respectively. The error bars depict the $0.9$-quantile of the estimated HF observation noise distribution. The plots appearing in the same row show that, for increasing noise and same EDs, both the CIs and the PIs become wider. This indicates that both the uncertainty about the underlying HF model and an unseen noise-contaminated HF observation increase. Moreover, the plots in the same columns reveal that, when the noise remains the same but the HF training data increases, the width of the CIs decreases. This means that our MFSM becomes more confident about where the regression model for the noise-free HF model lies. Also, the uncertainty about an unseen noise-contaminated HF observation, as shown by the PIs, decreases marginally. In this case, the PIs and CIs become more distinct, meaning that the uncertainty about a noise-contaminated HF observation is not anymore dominated by the regression model uncertainty.
@@
We observe that, as the HF ED increases, the CIs tend to converge to the regression model. This behavior aligns with our expectations, and reflects the fact that as more data becomes available, the epistemic uncertainty due to the lack of knowledge decreases and therefore, our MFSM exhibits increased confidence in its predictions. As regards the PIs, we notice that they tend to converge to a non-zero width, indicative of the amount noise in the HF observations. This behavior is also expected, as the noise in the HF observations arises from aleatory uncertainty, and is thus irreducible regardless of the amount of the available training data. Consequently, predictions for unseen observations will inherently carry this uncertainty.

### v2

@@ 2.4 Implementation
2.4.1 Construction of a multi-fidelity model
Starting from a set of high-fidelity experimental data $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$ and a lower-fidelity computational model ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$, we provide here the methodology to combine the two in a MFSM.
Our approach uses polynomial chaos expansion as a surrogate model in the hybrid correction scheme introduced in Equation 9, extending the works of Ng and Eldred (2012); Palar et al. (2016); Berchier (2016). The main motivations behind our choice of PCE as a surrogate in our MFSM methodology include its robustness to noise, its efficiency in terms of training, and its applicability in uncertainty quantification problems.
@@ 2.4.1 Construction of a multi-fidelity model
where $c_{\boldsymbol{\alpha}}\in\mathbb{R}$ are the coefficients of the multivariate polynomials $\Psi_{\boldsymbol{\alpha}},\,\boldsymbol{\alpha}\in{\mathcal{A}}$}. Each polynomial $\Psi_{\boldsymbol{\alpha}}$ is the product of univariate polynomials orthogonal with respect to the PDF $f_{\boldsymbol{X}_{i}}$ of the input variable $\boldsymbol{X}_{i}$, and characterized by the multi-index $\boldsymbol{\alpha}$. ${\mathcal{A}}\subset\mathbb{N}^{M}$ is the finite set of multi-indices of the polynomials, and it can be obtained from different truncation schemes, such as total-degree, low-rank or hyperbolic truncation (Marelli et al., 2022). The mean and variance of $\mathcal{M}(\boldsymbol{X})$ can be approximated by:
$\hat{\mu}_{\text{PC}}$ $=c_{\boldsymbol{0}}$ (15)
$\hat{\sigma}^{2}_{\text{PC}}$ $=\sum_{\begin{subarray}{c}\boldsymbol{\alpha}\in{\mathcal{A}}\\ \boldsymbol{\alpha}\neq\boldsymbol{0}\end{subarray}}c_{\boldsymbol{\alpha}}^{2}.$ (16)
For the calculation of the PCE coefficients $c_{\boldsymbol{\alpha}}$, we adopt a regression-based strategy, as exhaustively reviewed in Lüthen et al. (2021, 2022), because of its applicability to data-driven problems and robustness to noise (Torre et al., 2019). Specifically, we opt for a sparse regression solver, least angle regression (LARS) (Blatman and Sudret, 2011). Moreover, for the choice of the PCE basis we use degree adaptivity, as well as a total-degree truncation scheme for low-dimensional applications and hyperbolic truncation for higher-dimensional applications.
The first step in constructing our multi-fidelity surrogate model as in Equation 9 entails creating a surrogate of the low-fidelity model using PCE. For this purpose, we first sample $N_{\text{L}}$ realizations $\boldsymbol{{\mathcal{X}}}_{\text{L}}=\{\boldsymbol{x}_{\text{L}}^{(1)},...,\boldsymbol{x}_{\text{L}}^{(N_{\text{L}})}\}$ of the input random variables (e.g. through Latin hypercube sampling; McKay et al. (1979)) and obtain the corresponding model responses ${\mathcal{Y}}_{\text{L}}=\{y_{\text{L}}^{(1)},...,y_{\text{L}}^{(N_{\text{L}})}\}$. We then construct a PCE model
@@
as discussed above. By using ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$ in place of the original LF model ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$, we eliminate the need for the HF training set to be a subset of the LF one, as from now on we are able to obtain evaluations of ^$\widehat{{\mathcal{M}}}_{\text{L}}$ at a negligible cost. Moreover, we are able to remove any noise that may be present in the LF data in a general MFSM scenario. However, inaccuracies in the PCE surrogate of the LF model can impact the overall performance of the MFSM. Accurately approximating the LF model using PCE requires that the LF model response is sufficiently smooth. Moreover, highly non-linear LF models may require higher-order polynomials in the PCE to achieve accurate approximations. The use of degree-adaptive PCE allows for selecting the most appropriate degree, while sparse regression keeps the number of coefficients to be computed manageable and minimizes the risk of over-fitting. Nevertheless, in cases where the cost of evaluating the LF model is negligible, such as when the LF model is given by an analytical function, the LF model ${{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$ can be used directly in the MFSM. This eliminates the need to train ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$, thus avoiding any potential loss of accuracy.
@@
where $\mathcal{A}_{\text{L}}\cap\mathcal{A}_{\delta}$ is the set of multi-indices present in both the LF and the discrepancy expansions, $\mathcal{A}_{\text{L}}\text{\textbackslash}\mathcal{A}_{\delta}$ is the set of multi-indices present only in the LF expansion, and $\mathcal{A}_{\delta}\text{\textbackslash}\mathcal{A}_{\text{L}}$ the set of multi-indices present only in the discrepancy expansion. This last step is optional, and it is only possible when both ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$ and $\hat{\delta}(\boldsymbol{x})$ are PCEs. In general, different SMs could be used for either of these two models, in which case the combined expression in Equation 21 is not available. Thus, the corresponding step is omitted, and the multi-fidelity predictor is instead expressed as:
@@
According to the categorization by Peherstorfer et al. (2018), as discussed in the introduction of this paper, this multi-fidelity surrogate modeling method falls under the adaptation category, since the LF data is initially used to construct the LF surrogate model $\hat{\psi}_{\text{L}}$, which is then corrected through a discrepancy function and a scaling factor using HF model evaluations to predict the HF model.
In summary, the construction of the multi-fidelity surrogate model from a HF experimental data set $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$ and a LF computational model ${\mathcal{M}}_{\text{L}}(\boldsymbol{x})$ according to Equation 9 involves the following main steps:
@@
    Use the computed ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x}),\,\hat{\rho},\,\hat{\delta}(\boldsymbol{x})$ for the MF predictor as in Equation 22. Optionally, merge the LF and discrepancy expansions into one PCE.
In a broader multi-fidelity setting, a low-fidelity data set, or even a pre-trained surrogate of the LF model (not necessarily a PCE model) can be available instead of a LF computational model. In this situation, one can follow the same procedure to construct the MFSM, by simply omitting Step $1$, and in the second case also Step $2$. Furthermore, when a high-fidelity computational model is available instead of a HF data set, an additional step precedes Step $1$. This consists in using sampling to obtain a HF experimental design $(\boldsymbol{{\mathcal{X}}}_{\text{H}},{\mathcal{Y}}_{\text{H}})$. These adaptations accommodate the general multi-fidelity case, which need not strictly adhere to the framework of combining physical experiments and computer simulations.
@@
one can jointly estimate $\rho$ and the coefficients $c_{\alpha,\delta,\text{new}}$ using OLS. Another approach is based on alternating least squares (Chevreuil et al., 2015), where the estimates for $\rho$ and $\delta(\boldsymbol{x})$ are iteratively refined. The starting value for $\hat{\rho}$ in this joint optimization can be obtained from the 4th step of the algorithm described above.
If the LF model response is close to zero, the current estimator of $\rho$ given in Step 4 can be problematic. In these cases, one can employ an alternative estimator, computed as the ratio of the standard deviations of the HF and the LF models. These standard deviations can be computed, for example, according to Eq. 16 from PCEs trained on the available HF and LF data sets. Then,
$\hat{\rho}_{\text{alt}}=\frac{\hat{\sigma}_{\text{H}}}{\hat{\sigma}_{\text{L}}}.$ (24)
While this estimator of $\rho$ provides the additional benefit of being invariant to linear transforms between the HF and LF model responses, it can be sensitive to noise in the data. Therefore, one needs to use it with caution, and only in cases where minor noise is expected.
Finally, an alternative method that avoids these risks is similar to the first variation for joint estimation of $\rho$ and the coefficients $c_{\alpha,\delta}$ of the expansion of $\delta(\boldsymbol{x})$, but instead of performing a two-step joint estimation of $\rho$ and $c_{\alpha,\delta}$, it combines both estimations in a single step. Specifically, ^$\widehat{{\mathcal{M}}}_{\text{L}}(\boldsymbol{x})$ in Equation 22 is treated as a basis function alongside the basis functions ${\Psi}_{\alpha}\left(\boldsymbol{x}\right)$ of $\hat{\delta(\boldsymbol{x})}$. Then, a sparse regression technique, such as LARS or subspace pursuit (Dai and Milenkovic, 2009; Diaz et al., 2018) is employed to jointly estimate $\rho$ and $c_{\alpha,\delta}$, eliminating the need to fit each of them separately in an initial step.
In the applications discussed below, we opt for the method outlined in Steps 1-6 due to its simplicity. Notably, its performance closely paralleled that of the variations discussed above.
@@ 2.4.2 Construction of confidence and prediction intervals
One of the applications of the bootstrap estimator lies in constructing confidence intervals for regression models (Freedman, 1981). Bootstrap methods have been used to construct confidence and prediction intervals for regression model predictions in a number of single-fidelity studies (Heskes, 1996; Kumar and Srivistava, 2012). Moreover, while the application of bootstrap to provide local error estimates to PCE model predictions within a single-fidelity context has been previously studied (see Marelli and Sudret (2018)), its usage in the context of MFSM has not yet been explored.
@@
This means that a realization of $\varepsilon_{\text{H}}$ can be estimated as the difference between a HF observation and the bootstrap mean. Having $N_{\text{H}}$ noise samples $r^{(i)}$, we use maximum likelihood estimation (MLE) to infer the parameters of a zero-mean distribution. Within the scope of this work, we consider zero-mean variants of the classical Gaussian, Laplace, and Uniform distributions, but in the general case any zero-mean symmetrical or non-symmetrical distribution could be considered. Finally, we use the Bayesian information criterion (BIC; Schwartz (1978)) to select the most appropriate distribution among those considered.
@@ 3 Validation and results
In this section, the performance of the proposed framework for multi-fidelity surrogate modeling is illustrated on three examples of increasing complexity: an analytical example with one-dimensional input, a case study with a ten-dimensional finite-element model of a truss and its simply supported beam approximation, and a real-world application involving wind turbine simulations. In each application, the HF data contain noise, which is either naturally present (wind turbine application) or artificially introduced by us to replicate the real-world noise-contaminated setting (analytical 1-D example and truss model). The validation of the proposed framework comprises two parts: assessing the performance of our MFSM and appraising the confidence and prediction intervals, whose construction is described in the previous section.
@@ 3.1 MFSM performance and convergence evaluation
where $y_{\text{t}}^{(i)}$ equals the noise-free HF model response $\psi_{\text{H}}(\boldsymbol{x}_{\text{t}}^{(i)})$ at the test point $\boldsymbol{x}_{\text{t}}^{(i)}$, when this response is available (analytical 1-D example and truss model), or the noisy HF response $y_{\text{H}}(\boldsymbol{x}_{\text{t}}^{(i)})$ when the noise-free response is not known (wind turbine application), and $\mu_{y}$ is the mean value of the HF response.
@@ 3.3 Analytical 1-D example
where $x\sim\mathcal{U}\left[0,2\right]$. The HF and LF functions are depicted in Figure 2. For a data set of $1,000$ HF and LF samples, the Pearson correlation coefficient between the HF and LF data is 0.99, while the normalized root mean square error (NRMSE), normalized by the standard deviation of the HF model, is 0.18.
We artificially contaminate the HF data with additive noise that follows a Gaussian distribution $\varepsilon_{\text{H}}\sim\mathcal{N}(0,\,\sigma_{\varepsilon_{\text{H}}})$.
@@ 3.3.1 MFSM performance and convergence
Here, for each noise level, the HF ED varies from 10 to 50 data points, while the LF ED is fixed in all experiments to 100 data points. For the choice of basis in all PCEs involved in this example, we use degree adaptivity from degree 1 up to degree 15. For the computation of the validation errors shown in LABEL:fig:1-d_boxplot, we use $N_{test}=10^{5}$ noise-free HF data points generated using LHS in the input space.
@@
We observe that from as few as 10 HF training data points, our MFSM approach outperforms surrogate models trained solely on either HF or LF data. This distinction becomes particularly evident for lower noise levels within the HF data. Hence, when HF data is scarce and data of different fidelities is present, the value of employing MFSMs as opposed to single-fidelity SMs becomes apparent. We notice that when sufficient HF data is available, e.g., here, approximately 30 data points for $\varepsilon_{\text{H}}\sim{\mathcal{N}}(0,\,0.01\hat{\sigma}_{\text{H}})$, the MFSM and HF surrogate model performance is similar. In some instances, when a large amount of HF data is available, the MFSM performance can even be worse than that of the HF surrogate due to the bias introduced by the LF model. As discussed by Fernández-Godino et al. (2019), it is not guaranteed that MFSMs always outperform their single-fidelity counterparts. In addition, as the level of noise in the HF data increases, the difference in performance of the MFSM and the HF SM diminishes. Indeed, LABEL:fig:noise_4 shows a comparable performance between the two models, regardless of the size of the experimental design.
@@ 3.3.2 Confidence and prediction intervals
LABEL:fig:1d_CIs_PIs shows the $90\%$ CIs (blue area) and $90\%$ PIs (yellow area) for the MFSM prediction (blue line) in four different cases occurring from all combinations among a realization of higher/lower noise in the HF observations and a realization of a larger/smaller HF ED. Here, again the LF ED is fixed to 100 samples. The HF and LF training data in each case is visualised by the black error bars and the gray circles, respectively. The error bars depict the $0.9$-quantile of the estimated HF observation noise distribution. The plots appearing in the same row show that, for increasing noise and same EDs, both the CIs and the PIs become wider. This indicates that both the uncertainty about the underlying HF model and an unseen noise-contaminated HF observation increase. Moreover, the plots in the same columns reveal that, when the noise remains the same but the HF training data increases, the width of the CIs decreases. This means that our MFSM becomes more confident about where the noise-free HF model lies. Also, the uncertainty about an unseen noise-contaminated HF observation, as shown by the PIs, decreases marginally. In this case, the PIs and CIs become more distinct, meaning that the uncertainty about a noise-contaminated HF observation is not anymore dominated by the regression model uncertainty.
@@
We observe that, as the HF ED increases, the CIs tend to converge to the multi-fidelity regression model. This behavior aligns with our expectations, and reflects the fact that as more data becomes available, the epistemic uncertainty due to the lack of knowledge decreases and therefore, our MFSM exhibits increased confidence in its predictions. As regards the PIs, we notice that they tend to converge to a non-zero width, indicative of the amount of noise in the HF observations. This behavior is also expected, as the noise in the HF observations arises from aleatory uncertainty, and is thus irreducible regardless of the amount of the available training data. Consequently, predictions for unseen observations will inherently carry this uncertainty.